OpenStudy (anonymous):

Differentiate the function f(x)=log10(x^7 + 4)

OpenStudy (luigi0210):

Do you know what \(\frac{d}{dx}~log_{10} x\) is?

OpenStudy (anonymous):

yeah i guess, is it (7x^6)/(x^7 +4)log(10) ??

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

I would write it as \[\LARGE \frac{d}{dx}=\frac{7x^6}{(x^7+4)ln~10}\] but yea

OpenStudy (anonymous):

yeah i wrote it like that as my answer but it says its wrong :(